# Multi-group Uncertainty Quantification for Long-form Text Generation

### v1

@@ 3.2 Linear regressor algorithms
Next, we consider a set of algorithms that instead solve some optimziation problem for the purpose of calibration and conformal prediction. In these cases, one can naturally make them multi-group/valid by including group-membership (i.e., $g(X)=1$ for all $g\in\mathcal{G}$) in the optimization problem itself. Formally, we describe these linear regression based methods in Algorithms 4 and 5. Presented in this way, the methods for calibration vs. conformal prediction is reduced to a choice of loss function $L$. Again, we assume one has access to some calibration set for which one solves the optimization problem on.
@@
In the case of calibration, one can choose $L$ to be binary cross-entropy loss. In doing so, Algorithm 4 then describes Platt Scaling (PS) (PLATT, 1999), which can be described as fitting a logistic regression model to some set of model outputs to obtain calibrated probability scores. Algorithm 5 describes the multicalibrated version of Platt Scaling (Gopalan et al., 2022). Going forward, we refer to this method as Group Conditional Unbiased Logistic Regression (GCULR).
@@
For conformal prediction, we instead choose $L$ to be pinball loss. We refer to the non-group version of this method (Algorithm 4) as Conformalized Quantile Regression (CQR) (Romano et al., 2019), in which given some target coverage $1-\alpha$, we fit a linear quantile regression model that minimizes pinball loss.
In our conformal prediction setting, as described in Section 2.2, $X$ is an entire biography, or set of independent claims. Thus, to adapt quantile regression to long-form generation, we propose setting $f(X)$ to be a vector of uncertainty scores for each claim $x\in X$ [^4] Since each biography generated by the LLM may have a different number of claims, we apply interpolation to (un)squeeze the set of scores to a list of $K=25$ scores. Like in split conformal, the target is then the minimum threshold $r(X,Y)$ for which all claims above it are correct. In the multivalid case, we then add group features $g(X)$ to the optimization problem. A version of Algorithm 5 was first presented by Jung et al. (2022), and going forward, we will refer to this method as Group Conditional Conformalized Quantile Regression (GCCQR). Again, we set $f(X)$ be some vector of (interpolated) scores.
@@
1:  Input: scoring function $f$, loss function $L$
@@
1:  Input: scoring function $f$, loss function $L$, set of groups $\mathcal{G}$
@@
4 Empirical evaluation
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.32291 0.00875 0.00038 0.00022 0.00015*
@@ 4 Empirical evaluation
Mistral 7B Instruct marginal 0.30706 0.01163 0.00026 0.00029 0.00013*
@@
Table 1: We generate biographies for entities from Bio-NQ. We evaluate using ASCE, max gASCE, and average gASCE ($\downarrow$ better), comparing each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better-performing method for each pairing. * denotes the best-performing method across all methods evaluated.
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.463 0.218 0.190 0.196 0.181*
group max 0.574 0.293 0.260 0.267 0.247*
group mean 0.470 0.208 0.179 0.186 0.171*
Mistral 7B Instruct marginal 0.455 0.227 0.194 0.201 0.187*
group max 0.609 0.302 0.258 0.271 0.245*
group mean 0.466 0.216 0.180 0.190 0.174*
@@
Table 2: We generate biographies for entities from Bio-NQ. We evaluate using Brier score ($\downarrow$ better) marginally across the entire dataset, as well as within each subgroup (in terms of max and mean overall all groups). We compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better-performing method for each pairing. * denotes the best-performing method across all methods evaluated.
We focus our empirical evaluation on the problem of biography generation. Following Min et al. (2023a), we use an LLM to automate the process of decomposing biographies into individual claims and evaluating for factuality. Details can be found in Appendix A.
4.1 Dataset
@@
We evaluate on a large set of biographies by extracting 8,541 entities from Natural Questions dataset (Kwiatkowski et al., 2019), which consists of real queries issued to the Google search engine. We denote this dataset as Bio-NQ. Our motivation for choosing Natural Questions is that these extracted human entities should serve as a representative sample of public figures that users may prompt an LLM to generate information for. To form our dataset, for each question, we select all entities in either the question’s short answer or accompanying Wikipedia article. We then attempt to match them to their corresponding Wikidata entry. If a match exists and its Wikidata page’s property, if instance of, is equal to the value, human, we add the entity to our dataset.
4.2 Collecting group features

### v2

@@ 3.2 Linear regressor algorithms
Next, we consider algorithms that instead solve an optimization problem for the purpose of calibration and conformal prediction. In these cases, one can naturally make them multi-group/valid by including group-membership (i.e., $g(X)=1$ for all $g\in\mathcal{G}$) in the optimization problem itself. Formally, we describe these linear regression based methods in Algorithms 4 and 5. Presented in this way, the methods for calibration vs. conformal prediction is reduced to a choice of loss function $L$. Again, we assume one has access to some calibration set for which one solves the optimization problem on.
@@
For calibration, one can choose $L$ to be binary cross-entropy loss. In doing so, Algorithm 4 then describes Platt Scaling (PS) [Platt, 1999], which can be described as fitting a logistic regression model to some set of model outputs to obtain calibrated probability scores. [^5] A related calibration method to Platt scaling (PS) is temperature scaling (TS) [Guo et al., 2017], which was originally introduced for calibrating neural networks for multiclass classification and has been incorporated in work on calibrating NLP models [Sicilia et al., 2024]. We note, however, that in the binary classification setting (e.g., our setting where we identify if an output is correct or not), TS is mathematically equivalent to PS when there is no bias term and the weight takes on the form $\frac{1}{\tau}$, where $\tau$ is the temperature learned in TS. Algorithm 5 describes the multi-calibrated version of Platt Scaling. While not explicitly derived in their work, this multicalibration formulation can be traced back to Gopalan et al. [2022], who establish a hierarchy of notions for multicalibration and analyze multicalibration on functions trained with linear loss. Going forward, we refer to this method as Group Conditional Unbiased Logistic Regression (GCULR).
@@
For conformal prediction, we instead choose $L$ to be pinball loss. We refer to the non-group version of this method (Algorithm 4) as Conformalized Quantile Regression (CQR) [Romano et al., 2019], in which given some target coverage $1-\alpha$, we fit a linear quantile regression model that minimizes pinball loss.
In our conformal prediction setting, as described in Section 2.2, $X$ is an entire biography, or set of independent claims. Thus, to adapt quantile regression to long-form generation, we propose setting $f(X)$ to be a vector of uncertainty scores for each claim $x\in X$. Like in split conformal, the target is then the minimum threshold $r(X,Y)$ for which all claims above it are correct. In the multivalid case, we then add group features $g(X)$ to the optimization problem. A version of Algorithm 5 was first presented by Jung et al. [2022], and going forward, we will refer to this method as Group Conditional Conformalized Quantile Regression (GCCQR).
We note that in our experiments, each biography generated by the LLM may have a different number of claims, a setting in which prior work on conformal quantile regression does not account for. Consequently, we propose using interpolation to (un)squeeze the set of scores to a vector $f(X)$ of fixed size ($K=25$ in our experiments). While Mohri and Hashimoto [2024] only show that split conformal can be applied to this type of setting, our experiments demonstrate that quantile regression methods achieve similar performance for marginal (CQR) and multigroup (GCCQR) methods (Section 5).
@@
1:  Input: data distribution $\mathcal{D}$, scoring function $f$, loss function $L$
@@
1:  Input: data distribution $\mathcal{D}$, scoring function $f$, loss function $L$, set of groups $\mathcal{G}$
@@
Model Base Score Metric Uncalibrated HB IGHB PS GCULR
Llama self-consistency marginal 0.32291 0.00875 0.00038 0.00022 0.00015*
@@
P(True) marginal 0.11768 0.00451 0.00021* 0.00036 0.00022
group max 0.20701 0.06988 0.01798 0.06025 0.00697*
group mean 0.12682 0.01176 0.00328 0.00654 0.00145*
verb. conf. marginal 0.01642 0.00014 0.00055 0.00013* 0.00023
group max 0.06645 0.06634 0.01315 0.06709 0.00730*
group mean 0.02447 0.00738 0.00357 0.00750 0.00154*
Mistral self-consistency marginal 0.30706 0.01163 0.00026 0.00029 0.00013*
@@
P(True) marginal 0.06291 0.00074 0.00031 0.00047 0.00015*
group max 0.12293 0.06942 0.01417 0.07054 0.00587*
group mean 0.07173 0.00896 0.00309 0.00886 0.00132*
verb. conf. marginal 0.22229 0.00047 0.00036 0.00034 0.00015*
group max 0.33763 0.06922 0.01453 0.07080 0.00653*
group mean 0.23230 0.00869 0.00315 0.00878 0.00128*
@@
Table 1: We generate biographies using Llama 2 7B Chat and Mistral 7B Instruct for entities from Bio-NQ and compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR) on ASCE, max gASCE, and average gASCE ($\downarrow$ better). We test each method using the base scores: self-consistency, P(True), and verbalized confidence. We bold the better-performing method for each pairing and use * to denote the best-performing method across all methods.
Model Base Score Metric Uncalibrated HB IGHB PS GCULR
Llama self-consistency marginal 0.475 0.169 0.148 0.152 0.143*
group max 0.535 0.323 0.247 0.285 0.235*
group mean 0.479 0.169 0.148 0.152 0.143*
P(True) marginal 0.274 0.165 0.152 0.157 0.149*
group max 0.341 0.315 0.261 0.305 0.250*
group mean 0.277 0.165 0.152 0.157 0.148*
verb. conf. marginal 0.177 0.161 0.152 0.161 0.150*
group max 0.270 0.311 0.253 0.311 0.248*
group mean 0.177 0.160 0.152 0.160 0.149*
Mistral self-consistency marginal 0.471 0.186 0.164 0.159 0.152*
group max 0.554 0.333 0.285 0.250 0.235*
group mean 0.477 0.186 0.164 0.158 0.152*
P(True) marginal 0.237 0.175 0.164 0.174 0.161*
group max 0.304 0.318 0.259 0.317 0.249*
group mean 0.237 0.175 0.164 0.174 0.160*
verb. conf. marginal 0.397 0.175 0.164 0.175 0.161*
group max 0.427 0.318 0.259 0.318 0.249*
group mean 0.398 0.175 0.164 0.174 0.160*
@@
Table 2: We generate biographies using Llama 2 7B Chat and Mistral 7B Instruct for entities from Bio-NQ and compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR) on Brier score ($\downarrow$ better) marginally across the entire dataset, as well as within each subgroup (in terms of max and mean over all groups). We test each method using the base scores: self-consistency, P(True), and verbalized confidence. We bold the better-performing method for each pairing and use * to denote the best-performing method across all methods.
4 Empirical evaluation
We focus our empirical evaluation on the problem of biography generation, which we contend serves as a very suitable testbed for evaluating factuality and has been used as a benchmark in a variety of works in recent years. Outputting biographies offers one the ability to evaluate not only a set of objective and specific claims but also on a wide range of topics, which in turn allows us to explore a rich set of group functions for each person. Moreover, bias within biography generation has long been a studied issue, further motivating the problem of ensuring group-conditional guarantees. Like Min et al. [2023a], we use a language model to automate the process of decomposing biographies into claims and evaluating for factuality (Appendix B.1).
Dataset
@@ 4 Empirical evaluation
We evaluate on a large set of biographies by extracting 8,541 entities from the Natural Questions dataset [Kwiatkowski et al., 2019], which consists of real queries issued to the Google search engine. We denote this dataset as Bio-NQ. Our motivation for choosing Natural Questions is that these extracted human entities should serve as a representative sample of public figures that users may prompt an LLM about. For each question, we select all entities in either the question’s short answer or accompanying Wikipedia article. We then attempt to match them to their corresponding Wikidata entry. If a match exists and its Wikidata page’s property, if instance of, is equal to the value, human, we add the entity to our dataset Bio-NQ.
Collecting group features